# Beyond Markovian: Reflective Exploration via Bayes-Adaptive RL for LLM Reasoning

### v1

@@ 7.1 Experiment Setups
In addition to the synthetic experiment in Section 6, we evaluate BARL on LLM math reasoning tasks. Training is conducted on the Big-Math dataset (albalak2025big, ) across various models, including Qwen2.5-Math-1.5B, Qwen2.5-Math-7B (yang2024qwen25mathtechnicalreportmathematical, ), and DeepSeek-R1-Distill-Llama-8B (guo2025deepseek, ). Evaluation is performed on four benchmarks: GSM8K (cobbe2021training, ), MATH (hendrycks2021measuring, ), CollegeMath (tang2024mathscale, ), and OlympiadBench (he2024olympiadbench, ). During training, the maximum prompt length is set to $512$ and the maximum response length is set to $1024$. We provide the AIME and AMC results in Appendix B.1 and exclude them from our main evaluation due to their substantially longer context requirements, e.g., R1-Distill-Llama-8B has average response lengths of $2008$ and $1886$ tokens on AIME 2024 and AMC 2023, respectively. For BARL, we set $\beta=1$ and $|\mathcal{M}|=5$.
We compare BARL against two Markovian RL baselines that span outcome-reward and process-reward RL. For the outcome-reward GRPO baseline, we set its group size to $5$ for a fair comparison with BARL and, after performing a grid search over the KL-divergence coefficients $[0,0.001,0.005,0.01]$, adopt $0.005$ as it yields the best overall performance across all benchmarks. For the process-reward baseline, we adapt a variant of MRT (qu2025optimizing, ) by integrating the progress reward defined in (3.1) into the outcome reward, which we refer to as progress in the following sections. For all algorithms, we set the training and rollout batch sizes to $128$ and $1024$, respectively. We train the Qwen and Llama models for $110$ and $60$ iterations, respectively, defined w.r.t. the rollout batches. The temperature during online sampling is $1.0$ and is $0.0$ during evaluation. For both BARL and the progress baseline, we set the number of tokens for each reasoning step as $128$. All algorithms are implemented using the OpenRLHF codebase (hu2024openrlhf, ).
@@ 7.2 Experiment Results
We report the pass@1 accuracies in Table 1. All models are trained using three random seeds, and we calculate the mean and standard deviation of the resulting accuracies. For each algorithm, we report the results of the checkpoint that has the highest average accuracy.
Model GSM8K MATH CollegeMath OlympiadBench Average
Qwen2.5-Math-1.5B $40.0$ $34.1$ $6.6$ $21.8$ $25.6$
   GRPO $83.9\scriptstyle(\pm 0.5)$ $71.5\scriptstyle(\pm 0.4)$ $45.1\scriptstyle(\pm 0.3)$ $33.4\scriptstyle(\pm 0.4)$ $58.4\scriptstyle(\pm 0.3)$
   Progress $84.8\scriptstyle(\pm 0.6)$ $72.3\scriptstyle(\pm 0.4)$ $45.9\scriptstyle(\pm 0.3)$ $35.6\scriptstyle(\pm 0.2)$ $59.6\scriptstyle(\pm 0.2)$
   BARL $\textbf{86.0}\scriptstyle(\pm 0.6)$ $\textbf{72.7}\scriptstyle(\pm 0.3)$ $\textbf{46.8}\scriptstyle(\pm 0.2)$ $\textbf{35.8}\scriptstyle(\pm 0.4)$ $\textbf{60.3}\scriptstyle(\pm 0.1)$
Qwen2.5-Math-7B 59.1 53.7 21.9 19.0 38.4
   GRPO $90.3\scriptstyle(\pm 0.1)$ $77.6\scriptstyle(\pm 0.4)$ $47.0\scriptstyle(\pm 0.3)$ $39.0\scriptstyle(\pm 0.2)$ $63.5\scriptstyle(\pm 0.2)$
   Progress $91.1\scriptstyle(\pm 0.3)$ $78.8\scriptstyle(\pm 0.1)$ $47.2\scriptstyle(\pm 0.3)$ $41.1\scriptstyle(\pm 0.2)$ $64.5\scriptstyle(\pm 0.2)$
   BARL $\textbf{91.7}\scriptstyle(\pm 0.2)$ $\textbf{79.2}\scriptstyle(\pm 0.3)$ $\textbf{47.5}\scriptstyle(\pm 0.2)$ $\textbf{42.0}\scriptstyle(\pm 0.4)$ $\textbf{65.1}\scriptstyle(\pm 0.3)$
R1-Distill-Llama-8B $82.0$ $65.7$ $35.8$ $26.5$ $52.5$
   GRPO $85.7\scriptstyle(\pm 0.5)$ $\textbf{74.3}\scriptstyle(\pm 0.4)$ $39.6\scriptstyle(\pm 0.3)$ $36.0\scriptstyle(\pm 0.5)$ $58.9\scriptstyle(\pm 0.4)$
   Progress $\textbf{85.9}\scriptstyle(\pm 0.4)$ $73.9\scriptstyle(\pm 0.4)$ $39.8\scriptstyle(\pm 0.5)$ $35.3\scriptstyle(\pm 0.2)$ $58.7\scriptstyle(\pm 0.4)$
   BARL $85.4\scriptstyle(\pm 0.6)$ $73.9\scriptstyle(\pm 0.6)$ $\textbf{40.4}\scriptstyle(\pm 0.3)$ $\textbf{37.2}\scriptstyle(\pm 0.5)$ $\textbf{59.3}\scriptstyle(\pm 0.4)$
Table 1: Mean and standard deviation of the accuracies over three independent training runs.
It can be observed that BARL achieves higher accuracies across most benchmarks and models. It consistently outperforms Markovian RL baselines in terms of average accuracy, with larger gains observed on challenging benchmarks that demand effective exploration, such as CollegeMath and OlympiadBench. Notably, BARL delivers these gains with minimal computational overhead, which comes from calculating the probabilities of candidate answers at the end of each step, reusing the cache from CoT generations.
@@ 7.3 Ablation Studies
To qualitatively assess the improved token efficiency of BARL, we analyze the frequency of reflective behaviors across problems of varying difficulty levels, as shown in Figure 6. For each problem, we sample $6$ responses per model and define its difficulty level by the number of incorrect responses. We use keyword-based detections (liu2025understanding, ; yeo2025demystifying, ) to identify whether self-reflections appear in a response, and a problem is considered to exhibit self-reflection if at least one of its responses is identified. It can be observed that both models display fewer reflections on easier problems, and the base model exhibits a higher frequency of reflections despite achieving lower accuracies. This result reveals the weak correlation between the performance of LLMs and the response length or the frequency of reflections. Rather, the effectiveness of thinking tokens and the efficiency of explorations are the determining factors, which we study in the following ablation.
@@
Key Experiment Findings BARL consistently outperforms Markovian RL baselines with superior token efficiency. Performance correlates with the effectiveness of reflective explorations, rather than their frequency. Optimality in Markovian RL can be attained by policies that memorize training solutions yet fail to generalize, with no guarantees on the emergence of self-reflections.
@@ 8 Conclusion
Large Language Models (LLMs) trained via Reinforcement Learning (RL) have exhibited emergent behaviors such as self-reflective reasoning. Yet in conventional Markovian RL, exploration is confined to the training phase to identify action sequences that maximize cumulative reward, and resorts to pure exploitation at test time. Besides, the Markov assumption indicates the dependency on history only through the state. Thus, Markovian RL neither ensures the emergence of reflective exploration nor explains its benefits during testing. We propose to fill this gap with Bayes-Adaptive RL, which explicitly considers test-time performance by maximizing the expected return under a posterior of MDPs. Within this framework, we propose BARL, a novel algorithm for LLM reasoning that provides principled guidance for when and how to engage in reflective exploration. BARL enables efficient exploration through hypothesis elimination and strategy switching. Our experiments are conducted on both synthetic and mathematical reasoning tasks, where we show that BARL outperforms Markovian RL algorithms at test time and its exploration is more efficient. As future work, we plan to extend our approach to broader domains, such as coding and agentic tasks.
@@ Acknowledgements
Appendix A Proof of Theorem 4.2
@@
Consider a full binary tree of depth $T^{*}$, with leaf set $\mathcal{L}$ of size $|\mathcal{L}|=2^{T^{*}}$. The states are the tree nodes, with the initial state fixed as the root node. The actions include the moves from the parent nodes to the child nodes as well as a resetting action from the leaf node to the root node. The rewards are not known and differ in different MDP hypotheses. Specifically, the reward is defined as $r_{\mathcal{M}_{i}}(s)=\mathbbm{1}(s=s_{i})$, where $s_{i}$ is a unique leaf node for $\mathcal{M}_{i}$. The prior of MDPs is $p(\mathcal{M}_{i})=1/|\mathcal{L}|$, where $i=1,\cdots,|\mathcal{L}|$. The cumulative return is undiscounted with the minimum traverse steps as the horizon $T$. The episode terminates once the agent receives a $1$ reward.
@@
For any Markovian policy $\pi$, define $f(s)=p(\exists t\geq 0:s_{t}=s\mid\pi)$ and denote $p_{s}$ as the probability that $\pi$ goes left at an internal node $s$. For any left and right children $s_{L}$ and $s_{R}$ of node $s$, it holds that $f(s_{L})=f(s)p_{s}$ and $f(s_{R})=f(s)(1-p_{s})$. Thus, $f(s_{L})+f(s_{R})=f(s)$. By an induction on depth, it follows that at every depth $d$, $\sum_{s:\text{depth}(s)=d}f(s)=1$. In particular, $\sum_{l\in\mathcal{L}}f(l)=1$. Since the total return under $\mathcal{M}_{i}$ is $V(\pi\mid\mathcal{M}_{i})=p(\pi\text{ ends at leaf }l_{i})=f(l_{i})$, the expected return for the optimal Markovian policy is
@@
Consider the following deterministic Bayes-adaptive policy. At the root node, the policy picks any leaf $l$ with positive posterior, i.e., $p(\mathcal{M}_{l}\mid h_{t})>0$, then follows the unique shortest path to $l$. Two possible outcomes can occur: if the ground-truth reward $r(l)=1$, then the episodes terminates with the collected reward; if $r(l)=0$, then the agent eliminates the hypothesis $\mathcal{M}_{l}$ from the posterior by setting $p(\mathcal{M}_{l}\mid h_{t:t+T^{*}})=0$, and returns to the root to repeat the process on the remaining leaves. By construction, the expected return of this Bayes-Adaptive policy is $1$, which is an exponential improvement in $T^{*}$ over the $1/2^{T^{*}}$ return of the optimal Markovian policy. ∎

### v2

@@ 7.1 Experiment Setups
In addition to the synthetic experiment, we evaluate BARL on math tasks. We implement BARL across various models, including Qwen2.5-Math-1.5B, Qwen2.5-Math-7B (yang2024qwen25mathtechnicalreportmathematical), and DeepSeek-R1-Distill-Llama-8B (guo2025deepseek). Training is conducted on the Big-Math dataset (albalak2025big), and evaluation is performed on four benchmarks: GSM8K (cobbe2021training), MATH (hendrycks2021measuring), CollegeMath (tang2024mathscale), and OlympiadBench (he2024olympiadbench). During training, the maximum prompt length is set to $512$ and the maximum response length is set to $1024$. For BARL, we set $\beta=1$ and $|\mathcal{M}|=5$ in Algorithm 1.
We compare BARL against two RL baselines that span outcome- and process-reward RL. For the outcome-reward GRPO baseline, we set its group size to $5$ for a fair comparison with BARL and, after performing a grid search over the KL‐divergence coefficients $[0,0.001,0.005,0.01]$, adopt $0.005$ as it yields the best overall performance across all benchmarks. For the process-reward baseline, we adapt a variant of MRT (qu2025optimizing) by integrating the progress reward defined in (3.2) into the outcome reward, which we refer to as progress in the following sections. For all algorithms, we set the training and rollout batch sizes to $128$ and $1024$, respectively. We train the Qwen and Llama models for $110$ and $60$ iterations, respectively, defined w.r.t. the rollout batches. The temperature during online sampling is $1.0$ and is $0.0$ during evaluation. For both BARL and the progress baseline, we set the number of tokens for each reasoning step as $128$.
@@ 7.2 Experiment Results
We report the pass@1 accuracies in Table 1. All models are trained using three random seeds, and we calculate the mean and standard error of the resulting accuracies. For each algorithm, we report the results of the checkpoint that has the overall best performance on all benchmarks.
Model GSM8K MATH College Olympiad AIME AMC Average
Qwen-1.5B $40.0$ $34.1$ $6.6$ $21.8$ 16.7 32.5 $25.3$
     GRPO $83.9\scriptstyle(\pm 0.5)$ $71.5\scriptstyle(\pm 0.4)$ $45.1\scriptstyle(\pm 0.3)$ $31.6\scriptstyle(\pm 0.4)$ $\textbf{17.8}\scriptstyle(\pm 0.9)$ $55.8\scriptstyle(\pm 0.7)$ $51.0\scriptstyle(\pm 0.3)$
     Progress $84.8\scriptstyle(\pm 0.6)$ $72.1\scriptstyle(\pm 0.4)$ $45.9\scriptstyle(\pm 0.3)$ $35.5\scriptstyle(\pm 0.3)$ $14.4\scriptstyle(\pm 0.9)$ $55.8\scriptstyle(\pm 0.7)$ $51.4\scriptstyle(\pm 0.2)$
     BARL $\textbf{85.8}\scriptstyle(\pm 0.1)$ $\textbf{72.7}\scriptstyle(\pm 0.3)$ $\textbf{46.8}\scriptstyle(\pm 0.2)$ $\textbf{35.8}\scriptstyle(\pm 0.3)$ $\textbf{17.8}\scriptstyle(\pm 0.9)$ $\textbf{60.8}\scriptstyle(\pm 0.7)$ $\textbf{53.3}\scriptstyle(\pm 0.4)$
Qwen-7B 59.1 53.7 21.9 19.0 13.3 42.5 34.9
     GRPO $90.3\scriptstyle(\pm 0.1)$ $77.6\scriptstyle(\pm 0.4)$ $47.0\scriptstyle(\pm 0.3)$ $38.5\scriptstyle(\pm 0.2)$ $24.5\scriptstyle(\pm 0.4)$ $65.0\scriptstyle(\pm 1.2)$ $57.1\scriptstyle(\pm 0.2)$
     Progress $91.1\scriptstyle(\pm 0.3)$ $78.8\scriptstyle(\pm 0.1)$ $46.7\scriptstyle(\pm 0.1)$ $41.1\scriptstyle(\pm 0.2)$ $24.5\scriptstyle(\pm 0.4)$ $65.4\scriptstyle(\pm 1.4)$ $57.9\scriptstyle(\pm 0.2)$
     BARL $\textbf{91.7}\scriptstyle(\pm 0.2)$ $\textbf{79.2}\scriptstyle(\pm 0.3)$ $\textbf{47.5}\scriptstyle(\pm 0.2)$ $\textbf{42.0}\scriptstyle(\pm 0.4)$ $\textbf{29.0}\scriptstyle(\pm 0.4)$ $\textbf{66.7}\scriptstyle(\pm 1.4)$ $\textbf{59.4}\scriptstyle(\pm 0.3)$
Llama-8B $82.0$ $65.7$ $35.8$ $26.5$ $10.0$ $42.5$ $43.7$
     GRPO $85.7\scriptstyle(\pm 0.5)$ $\textbf{74.3}\scriptstyle(\pm 0.4)$ $39.6\scriptstyle(\pm 0.3)$ $36.0\scriptstyle(\pm 0.5)$ $16.7\scriptstyle(\pm 0.8)$ $60.4\scriptstyle(\pm 0.3)$ $52.1\scriptstyle(\pm 0.4)$
     Progress $\textbf{85.9}\scriptstyle(\pm 0.4)$ $73.9\scriptstyle(\pm 0.4)$ $39.8\scriptstyle(\pm 0.5)$ $35.3\scriptstyle(\pm 0.2)$ $15.0\scriptstyle(\pm 0.8)$ $55.8\scriptstyle(\pm 0.7)$ $51.0\scriptstyle(\pm 0.4)$
     BARL $85.4\scriptstyle(\pm 0.6)$ $73.9\scriptstyle(\pm 0.6)$ $\textbf{40.4}\scriptstyle(\pm 0.3)$ $\textbf{37.2}\scriptstyle(\pm 0.5)$ $\textbf{17.8}\scriptstyle(\pm 0.9)$ $\textbf{61.7}\scriptstyle(\pm 0.7)$ $\textbf{52.7}\scriptstyle(\pm 0.4)$
Table 1: Mean and standard error of the accuracies over three independent training runs.
We observe that BARL achieves higher accuracies across most benchmarks and models. It consistently outperforms the two conventional RL baselines in terms of average accuracy, with the most significant gains observed on challenging benchmarks that demand effective exploration. Notably, BARL delivers these gains with minimal computational overhead, which comes from calculating the probabilities of candidate answers at the end of each step, reusing the prefix cache from CoT generations. We defer the curves of accuracies and response lengths to Appendix B.1 and B.2.
@@ 7.3 Ablation Studies
To qualitatively assess the improved token efficiency of BARL, we analyze the frequency of reflective behaviors across problems of varying difficulty levels, as shown in Figure 6. For each problem, we sample $6$ responses per model and define its difficulty level by the number of incorrect responses. We use keyword-based detections (liu2025understanding; yeo2025demystifying) to identify whether self-reflections appear in a response, and a problem is considered to exhibit self-reflection if at least one of its responses is identified. It can be observed that both models display fewer reflections on easier problems, and the base model exhibits a higher frequency of reflections despite achieving lower accuracies. This result reveals (1) the base model has acquired superficial stylistic reflection patterns during pre-training (e.g., from CoT annealing data) without effective exploration, whereas BARL yields more effective reflective exploration rather than stylistic patterns; and (2) the weak correlation between the performance of LLMs and the response length or the frequency of reflections. Rather, the effectiveness of thinking tokens and the efficiency of explorations are the determining factors, which we study in the following ablation.
@@
Key Experiment Findings BARL consistently outperforms conventional RL baselines with superior token efficiency. Performance correlates with the effectiveness of reflective explorations, rather than their frequency. Optimality in conventional RL can be attained by policies that memorize training solutions yet fail to generalize, with no guarantees on the emergence of self-reflections.
@@ 8 Conclusion
LLMs trained via RL have exhibited emergent behaviors such as self-reflections, which offer a potential way for test-time scaling by backtracking to a previous state and generating a different action distribution with updated context. However, conventional RL training neither ensures the emergence of reflective exploration nor explains its benefits. The optimal RL policies are typically Markovian and have no incentive to enrich a revisited state with additional context since they depend on the history only through the state. Instead, RL exploration is confined to the training phase to identify action sequences that maximize cumulative rewards, and resorts to pure exploitation when the policy parameters are frozen after deployment. We propose to ground reflective exploration with a Bayesian RL objective that maximizes expected deployment-time return under a posterior distribution over MDPs induced by the training data. To maximize this objective, we propose BARL, which provides step-level guidance for when and how to engage in reflective exploration. BARL enables efficient exploration through hypothesis elimination and strategy switching. Our experiments are conducted on both synthetic and mathematical reasoning tasks, where we show that BARL outperforms conventional RL algorithms for evaluation benchmarks, and its exploration is more efficient.
@@ Acknowledgements
Appendix A Proofs
A.1 Proof of Theorem 4.1
@@ Appendix A Proofs
Let $\pi^{*}$ be any optimal policy (possibly history-dependent). Suppose $\pi^{*}$ prescribes different action distributions at the same state $s$ depending on the history that led to $s$.
Let $\delta_{1}$ and $\delta_{2}$ be two such distributions with $\mathbb{E}_{a\sim\delta_{1}}Q^{*}(s,a)>\mathbb{E}_{a\sim\delta_{2}}Q^{*}(s,a)$, where $Q^{*}$ is the optimal value in RL. Then construct $\pi^{\prime}$ that behaves like $\pi^{*}$ everywhere except it always uses $\delta_{1}$ whenever the state is $s$ (ignoring history). The future trajectories from any visit to $s$ are stochastically identical under $\pi^{*}$ and $\pi^{\prime}$, except that $\pi^{\prime}$ gets strictly higher expected return from every time $s$ is visited after histories that would have triggered $\delta_{2}$ under $\pi^{*}$.
Thus, $V^{\pi^{\prime}}(s_{0})>V^{\pi^{*}}(s_{0})$ for some state, contradicting optimality of $\pi^{*}$. Therefore, this optimal policy must prescribe the same action distribution at state $s$ regardless of history. In other words, there exists an optimal Markovian policy. In the following, we will prove by contradiction that policies that exhibit reflective exploration cannot be Markovian.
Assume $\pi$ is a Markovian policy. Then there exists a mapping $\mu:\mathcal{S}\to\Delta(\mathcal{A})$ such that, for every history $h_{t}$, $\pi(\cdot\mid h_{t})=\mu(\cdot\mid s(h_{t}))$. Define the process $X_{t}:=s(h_{t})$ and set $\upsilon:=\mu$. By construction, $\pi(\cdot\mid h_{t})=\upsilon(\cdot\mid X_{t})$ for all $t$.
According to Definition 3.2, there is a measurable $\psi$ such that $X_{t}=\psi(\varphi(s_{t}))$ and thus $s(h_{t})=\psi(\varphi(s(h_{t})))$. Let $s,s^{\prime}$ be reachable states and $\varphi(s)=\varphi(s^{\prime})$. Then $s=\psi(\varphi(s))=\psi(\varphi(s^{\prime}))=s^{\prime}$.
Therefore, there exist $t_{1}>t_{2}$ and histories $h_{t_{1}},h_{t_{2}}$ such that
$\varphi(s(h_{t_{1}}))=\varphi(s(h_{t_{2}})),\qquad\pi(\cdot\mid h_{t_{1}})\neq\pi(\cdot\mid h_{t_{2}}).$
Thus, we obtain $s(h_{t_{1}})=s(h_{t_{2}})=:s_{*}$. But then, using the Markov property of $\pi$, we have $\pi(\cdot\mid h_{t_{1}})=\mu(\cdot\mid s_{*})=\pi(\cdot\mid h_{t_{2}})$, which contradicts $\pi(\cdot\mid h_{t_{1}})\neq\pi(\cdot\mid h_{t_{2}})$. Therefore, $\pi$ is not Markovian. ∎
A.2 Proof of Theorem 4.2
Before the proof, we first provide several useful definitions and lemmas.
Definition A.1 (Epistemic POMDP).
We define the epistemic POMDP $\mathcal{P}\coloneqq(\mathcal{Z},\mathcal{O},\mathcal{A},P,r,\rho)$ where the hidden state space as $\mathcal{Z}\coloneqq\mathcal{S}\times\mathcal{M}$, with state $z_{t}=(s_{t},M)$; action space $\mathcal{A}$; observation space $\mathcal{O}\coloneqq\mathcal{S}$ with an observation map $O((s,M))=s$; transition kernel $P\bigl((s_{t+1},M^{\prime})\mid(s_{t},M),a_{t}\bigr)=\mathbf{1}\{M^{\prime}=M\}\,P_{M}(s_{t+1}\mid s_{t},a_{t})$; reward $r((s,M),a)=r_{M}(s,a)$; and initial state distribution $\rho((s_{0},M))\coloneqq p(M\mid D)\,\rho_{M}(s_{0})$, where $\rho_{M}$ is the initial-state distribution in $M$. For a policy $\pi$, let $J_{\mathcal{P}}(\pi)$ denote its expected return in this POMDP.
Lemma A.2 (Equivalence of Objectives).
For any policy $\pi$, $\pi$ is Bayes-optimal if and only if it is optimal for the epistemic POMDP $\mathcal{P}$, i.e.,
$J_{\mathcal{P}}(\pi)\;=\;J_{\mathrm{Bayes}}(\pi).$
Proof.
By construction of the initial distribution,
$(s_{0},M)\sim\rho\quad\Longleftrightarrow\quad M\sim p(M\mid D),\;s_{0}\sim\rho_{M}.$
Conditioned on $M$, the state transitions and rewards in $\mathcal{P}$ coincide exactly with those of the MDP $M$ under policy $\pi$, and the observation $o_{t}=s_{t}$ carries no additional information beyond $s_{t}$. Thus,
$J_{\mathcal{P}}(\pi)=\mathbb{E}_{M\sim p(M\mid D)}[J_{M}(\pi)]=J_{\mathrm{Bayes}}(\pi),$
which proves the claim. Similar results are also presented in [singh1994learning, duff2002optimal]. ∎
Now we are ready to prove Theorem 4.2.
Proof of Theorem 4.2.
By Lemma A.2, Bayes-optimal policies coincide with optimal policies for the epistemic POMDP $\mathcal{P}$.
It is a standard result in POMDP theory that there exists an optimal policy that is a function only of the belief state $\beta_{t}$, where
$\beta_{t}(z)\coloneqq\mathbb{P}(z_{t}=z\mid h_{t},D).$
Thus there exists a policy $\pi^{\star}$ and a measurable mapping $\widetilde{\mu}^{\star}$ such that
$\pi^{\star}(a_{t}\mid h_{t})=\widetilde{\mu}^{\star}(a_{t}\mid\beta_{t}).$
In the epistemic POMDP, the hidden state is $z_{t}=(s_{t},M)$ and the observation is $o_{t}=s_{t}$. Therefore, the belief factors as
$\beta_{t}(s,M)=\mathbf{1}\{s=s_{t}\}\,p(M\mid h_{t},D),$
so $\beta_{t}$ is uniquely determined by the pair $(s_{t},p(\cdot|h_{t},D))$. Since the offline posterior $p(M|D)$ is fixed, the mapping $b(h_{t})(M)$ is in one-to-one correspondence with $p(M|h_{t},D)$, and hence with $\beta_{t}$. Thus there exists a measurable function $\mu^{\star}$ such that for all $h_{t}$,
$\widetilde{\mu}^{*}(a_{t}\mid\beta_{t})=\mu^{\star}(a_{t}\mid s_{t},b(h_{t})).$
Therefore, we obtain
$\pi^{\star}(a_{t}\mid h_{t})=\widetilde{\mu}^{\star}(a_{t}\mid\beta_{t})=\mu^{\star}(a_{t}\mid s_{t},b(h_{t})),$
so $\pi^{*}$ is uncertainty-adaptive. Since $\pi^{*}$ is optimal for $\mathcal{P}$, it is Bayes-optimal for $J_{\mathrm{Bayes}}$ by Lemma A.2, completing the proof. ∎
A.3 Proof of Theorem 4.3
Proof.
@@
Consider a full binary decision structure of depth $T^{*}$, with leaf set $\mathcal{L}$ of size $|\mathcal{L}|=2^{T^{*}}$. The states are the tree nodes, with the initial state fixed as the root node. The actions include the moves from the parent nodes to the child nodes as well as a resetting action from the leaf node to the root node. The rewards are not known and differ in different MDP hypotheses. Specifically, the reward is defined as $r_{\mathcal{M}_{i}}(s)=\mathbbm{1}(s=s_{i})$, where $s_{i}$ is a unique leaf node for $\mathcal{M}_{i}$. The prior of MDPs is $p(\mathcal{M}_{i})=1/|\mathcal{L}|$, where $i=1,\cdots,|\mathcal{L}|$. The cumulative return is undiscounted with the minimum traverse steps as the horizon $T$. The episode terminates once the agent receives a $1$ reward.
@@
For any Markovian policy $\pi$, define $f(s)=p(\exists t\geq 0:s_{t}=s\mid\pi)$. We define $p_{s}$ as the probability that $\pi$ goes left at an internal node $s$. By the Markov property, for any left and right children $s_{L}$ and $s_{R}$ of node $s$, it holds that $f(s_{L})=f(s)p_{s}$ and $f(s_{R})=f(s)(1-p_{s})$. Thus, $f(s_{L})+f(s_{R})=f(s)$. By a simple induction on depth, it follows that at every depth $d$, $\sum_{s:\text{depth}(s)=d}f(s)=1$. In particular, $\sum_{l\in\mathcal{L}}f(l)=1$. Since the total return under $\mathcal{M}_{i}$ is $V(\pi\mid\mathcal{M}_{i})=p(\pi\text{ ends at leaf }l_{i})=f(l_{i})$, the expected return for the optimal Markovian policy is
@@
We construct the uncertainty-adaptive policy as follows. Let $\mathcal{C}=\{\mathcal{M}_{1},\dots,\mathcal{M}_{|\mathcal{L}|}\}$ be the set of models that are still consistent with the observed rewards in $h_{t}$. When the agent visits a leaf $s$ and observes reward $r\in\{0,1\}$, we update $\mathcal{C}\leftarrow\{\mathcal{M}\in\mathcal{C}:r_{\mathcal{M}}(s)=r\}$. For each $i$, there is exactly one leaf $s$ with $r_{\mathcal{M}_{i}}(s)=1$. Hence, under any true MDP $\mathcal{M}^{*}$, after at most $|\mathcal{L}|$ visits to leaves we must have found the reward leaf, and obtain a return of $1$. ∎